# Assessing the Feasibility of Internet-Sourced Video for Automatic Cattle Lameness Detection

### v2

@@ II-C2 ConvLSTM2D
Figure 3: Confusion matrix of both models under the same test data.
II-D Model Evaluation
The four performance evaluation measures (accuracy, precision, recall, and F1-score) used in this study are associated with the confusion matrix; therefore, both are presented gradually. The values of all evaluation criteria range from 0% to 100%, where a higher value indicates stronger performance. In the mathematical framework for model evaluation, $n$ represents the total instances, $y_{i}$ denotes the true label for instance $i$, and $\hat{y}_{i}$ signifies the model’s predicted label. The indicator function $I(\text{condition})$ returns 1 if the condition is true, otherwise 0. Summation, denoted by $\sum_{i=1}^{n}$, calculates the cumulative value across all instances. These components enable quantitative assessment by comparing predicted ($\hat{y}_{i}$) and true ($y_{i}$) labels, facilitating calculation of the evaluation metrics, providing a comprehensive evaluation of model performance.
@@ II-D Model Evaluation
III Results
The dataset utilized in this study comprised 50 videos, with 50% depicting defective cattle and 50% depicting non-defective cattle. Of these 50 videos, 30 were used for model training, and 20 were used for validating the 3D CNN and ConvLSTM2D models. This project was developed using the Python programming language and implemented within the Google Colab Pro environment. Data preprocessing tasks were handled by OpenCV and MoviePy, supplemented by Keras’ ImageDataGenerator for data augmentation. Furthermore, TensorFlow, NumPy, Pandas, Scikit-learn, and Matplotlib were employed for model construction, training, performance evaluation, and visualization. The classification results achieved by the models were calculated and presented below.
The models generate predictions for each individual frame within each video in the test data. It is assumed that if the model correctly predicts the majority of frames in a video, then the entire video is considered correctly classified, despite the presence of a few incorrect frame-level predictions. The confusion matrices for both models are presented in Figure 3. As shown in Figures 3(a) and 3(b), a small number of misclassifications were observed, including 1 false positives and 1 false negative for 3D CNN, and 1 false positives and 2 false negatives for ConvLSTM2D.
The 3D CNN model demonstrated notable performance in classifying the cattle data. Table III indicates an accuracy of 90%, along with strong average precision, recall, and F1-scores. On the other hand, ConvLSTM2D achieved comparatively lower performance, with accuracy, precision, recall, and F1-score values of 85%, 90%, 81.82%, and 85.71%, respectively. Overall, the confusion matrices and classification reports highlight the models’ effectiveness in classifying healthy and defective cattle.
TABLE III: Lameness classification results of proposed 3D CNN and ConvLSTM2D
@@ III Results
IV Discussion
TABLE IV: Performance comparison of this study with existing cattle lameness classification models
@@ IV Discussion
Table IV presents a performance comparison between eight prior studies and the proposed model. The table evaluates models based on accuracy, precision, recall, and f1-score. The final row of the table displays the results of this study, which employed a 3D CNN model, while the preceding rows summarize the findings from other relevant studies that utilized various machine learning and deep learning classifiers. The model used in this study achieved a notable accuracy of 90%. More significantly, it demonstrated the highest precision among the studies where precision values were available, reaching 90.9%. This indicates that the model exhibits a strong ability to minimize false positives, correctly identifying positive cases with high confidence. Additionally, the 3D CNN model also yielded a robust F1-score of 90.91%. When compared to other approaches, the 3D CNN model in this study demonstrated competitive performance, whereas study [15] employed C3D-ConvLSTM achieved a slightly higher accuracy of 90.32% but a lower precision of 90.62%.
@@
IV-A Limitation and future plan

### v3

@@ 3.3.2 ConvLSTM2D
3.4 Model Evaluation
The four performance evaluation measures (accuracy, precision, recall, and F1-score) used in this study are associated with the confusion matrix; therefore, both are presented gradually. The values of all evaluation criteria range from 0% to 100%, where a higher value indicates stronger performance. In the mathematical framework (equation 1, 2, 3, and 4) for model evaluation, $n$ represents the total instances, $y_{i}$ denotes the true label for instance $i$, and $\hat{y}_{i}$ signifies the model’s predicted label. The indicator function $I(\text{condition})$ returns 1 if the condition is true, otherwise 0. Summation, denoted by $\sum_{i=1}^{n}$, calculates the cumulative value across all instances. These components enable quantitative assessment by comparing predicted ($\hat{y}_{i}$) and true ($y_{i}$) labels, facilitating calculation of the evaluation metrics, providing a comprehensive evaluation of model performance.
@@ 3.4 Model Evaluation
4 Results
This project was developed using the Python programming language and implemented within the Google Colab Pro environment. Data preprocessing tasks were handled by OpenCV and MoviePy, supplemented by Keras’ ImageDataGenerator for data augmentation. Furthermore, TensorFlow, NumPy, Pandas, Scikit-learn, and Matplotlib were employed for model construction, training, performance evaluation, and visualization.
(a) 3DCNN
(b) ConvLSTM2D
Figure 4: Confusion matrix of both models under the same test data.
The dataset utilized in this study comprised 50 videos, with 50% depicting defective cattle and 50% depicting non-defective cattle. Of these 50 videos, 30 were used for model training, and 20 were used for validating the 3D CNN and ConvLSTM2D models. In the test phase, 500 frames were selected from 20 test videos (10 normal and 10 lame) to evaluate the performance of both models individually. Frame sampling (25 frames per video) and frame size ($224\times 224$ pixels) were kept consistent with the training phase to ensure a fair assessment. Augmentation techniques were deliberately avoided for the test data to preserve the original video characteristics, enabling an unbiased evaluation on unaltered, real-world data. Each video was classified based on a majority vote from its 25 frame-level predictions; a video was considered correctly classified if the model accurately predicted the majority of its frames. Ultimately, the classification report was presented based on video-level performance. This method ensures a robust evaluation while accommodating the variable lengths and conditions present in real-world video data. The classification results achieved by the models were calculated and presented here.
Table 4: Lameness classification results of proposed 3D CNN and ConvLSTM2D
@@ 4 Results
The confusion matrices for both models are presented in Figure 4. As shown in Figures 4a and 4b, a small number of misclassifications were observed, including 1 false positives and 1 false negative for 3D CNN, and 1 false positives and 2 false negatives for ConvLSTM2D. The 3D CNN model demonstrated notable performance in classifying the cattle data. Table 4 indicates an accuracy of 90%, along with strong average precision, recall, and F1-scores. On the other hand, ConvLSTM2D achieved comparatively lower performance, with accuracy, precision, recall, and F1-score values of 85%, 90%, 81.82%, and 85.71%, respectively. Overall, the confusion matrices and classification reports highlight the models’ effectiveness in classifying healthy and defective cattle.
4.1 Qualitative Evaluation of Model Predictions
Along with the numerical evaluation, we also reviewed some of the test videos to understand where the models performed well or struggled. This qualitative analysis helps highlight practical strengths and weaknesses that may help to understand the model performance more clearly. Figure 5 presents representative frames from the test set illustrating the classification outcomes of both models. The top row corresponds to the 3D CNN model, while the bottom row shows the results of the ConvLSTM2D model. Each row includes one sample each of true positive (TP), true negative (TN), false positive (FP), and false negative (FN) predictions.
  1. 1.
    Correct Classifications: The 3D CNN model consistently produced correct predictions when gait abnormalities were visually prominent, for instance, limping or dragging a leg (TP, first column of Figure 5). Similarly, the second column shows a healthy cow walking with a balanced gait and straight backline, which the 3D CNN correctly classified as a TN. Notably, the same frame was misclassified by ConvLSTM2D, which labeled the healthy cow as lame. This inconsistency may indicate a limitation of the ConvLSTM2D model in handling slight movements in walking or staying reliable when appearances change.
  2. 2.
    Misclassifications: Though the overall results show strong performance, both models exhibited some misclassifications. The third and fourth columns of Figure 5 illustrate misclassification cases. The 3D CNN produced a FP in one instance, likely due to motion blur or transient posture changes.
  3. 3.
    Postural Differences in Lame vs. Healthy Cattle: In addition to prediction behavior, we observed distinct physical posture differences between lame and healthy cattle. As shown in Figure 2, the lame cow’s body forms a noticeable arc from the head to the tail. In this frames, the typical shoulder hump appears less defined, almost blending into the backline, possibly as a way to relieve pain or pressure while walking or standing. On the other side, the figure shows healthy cows are with a straight and level backline, an upright head, and a sharp, clearly visible shoulder hump. This posture indicates a normal body mechanics and balanced movement. These visual distinction offers meaningful insights into lameness and could support future model development by combining geometric body features with spatiotemporal motion data.
Overall, the qualitative examples in Figure 5 align with the trends seen in the confusion matrices (Figure 4) and classification scores (Table 4), and they further emphasize the ConvLSTM2D model’s relative weakness in cases with subtle or ambiguous visual patterns.
Figure 5: Sample frames from the test dataset illustrating classification outcomes of the two models. The upper row displays results from the 3D CNN model, and the lower row shows results from the ConvLSTM2D model. Each row includes examples of true positive (TP), true negative (TN), false positive (FP), and false negative (FN) predictions, respectively. These frames qualitatively highlight both correct classifications and misclassification scenarios.
5 Discussion
Table 5: Performance comparison of this study with existing cattle lameness classification models
@@ 5 Discussion
Table 5 presents a performance comparison between eight prior studies and the proposed model. The table evaluates models based on accuracy, precision, recall, and f1-score. The final row of the table displays the results of this study, which employed a 3D CNN model, while the preceding rows summarize the findings from other relevant studies that utilized various machine learning and deep learning classifiers. The model used in this study achieved a notable accuracy of 90%. More significantly, it demonstrated the highest precision among the studies where precision values were available, reaching 90.9%. This indicates that the model exhibits a strong ability to minimize false positives, correctly identifying positive cases with high confidence. Additionally, the 3D CNN model also yielded a robust F1-score of 90.91%. When compared to other approaches, the 3D CNN model in this study demonstrated competitive performance, whereas study [16] employed C3D-ConvLSTM achieved a slightly higher accuracy of 90.32% but a lower precision of 90.62%.
@@
This project’s computational aspects were implemented in Python using TensorFlow and Keras, and conducted in the Google Colab Pro environment equipped with an NVIDIA Tesla T4 GPU. The training set comprised 1,500 augmented frames from 30 videos, while the evaluation was performed on 500 frames from 20 test videos. The 3D CNN model demonstrated faster convergence and lower computational demand, with each training epoch requiring approximately 1 minutes. In contrast, the ConvLSTM2D model, which incorporates recurrent layers for temporal modeling, incurred slightly higher computational overhead, requiring about 1-2 minutes per epoch. Despite these differences, both models remained computationally feasible to train and evaluate using a cloud-based consumer-grade platform.
5.1 Limitation and future plan
@@
  • We did not benchmark recent transformer-based video models (e.g., TimeSformer, Video-Swin). Future work will explore.
@@ 6 Conclusions
Author Contributions: Md Fahimuzzman Sohan contributed to conceptualization, methodology, formal analysis, software implementation, and manuscript drafting. Raid Alzubi contributed to data curation, validation, and interpretation of results. Hadeel Alzoubi and Eid Albalawi contributed to methodology refinement and critical manuscript review. A. H. Abdul Hafez supervised the project, provided resources, and contributed to manuscript editing and final approval. All authors have read and approved the final manuscript.
Data availability The dataset used and analysed during the current study come from third parties that are publicly available. We commit to open-sourcing the codebase and datasets after the paper is accepted.
Declarations
Conflict of interest The authors declare no competing interests.
Ethics approval This study used publicly available video recordings of dairy cows from YouTube to develop and evaluate a spatiotemporal deep learning model for lameness detection. No animals were directly involved in experiments or handling, and all videos were collected in accordance with the platform’s terms of service. Ethical approval was not required as the study did not involve any live animal experimentation.